What volume of solution is needed to dissolve 35.6 grams of KOH into a 71% solution?



1
Expert's answer
2021-03-04T06:44:19-0500

m(total) = 35.6 / 0.71 = 50.1 (g)

V(water) = m(water) = 50.1 - 35.6 = 14.5 (ml)